# MetaMixer Is All You Need

### v1

@@ Abstract
Transformer, composed of self-attention and Feed-Forward Network (FFN), has revolutionized the landscape of network design across various vision tasks. While self-attention is extensively explored as a key factor in performance, FFN has received little attention. FFN is a versatile operator seamlessly integrated into nearly all AI models to effectively harness rich representations. Recent works also show that FFN functions like key-value memories. Thus, akin to the query-key-value mechanism within self-attention, FFN can be viewed as a memory network, where the input serves as query and the two projection weights operate as keys and values, respectively. Based on these observations, we hypothesize that the importance lies in query-key-value framework itself rather than in self-attention. To verify this, we propose converting self-attention into a more FFN-like efficient token mixer with only convolutions while retaining query-key-value framework, namely FFNification. Specifically, FFNification replaces query-key and attention coefficient-value interactions with large kernel convolutions and adopts GELU activation function instead of softmax. The derived token mixer, FFNified attention, serves as key-value memories for detecting locally distributed spatial patterns, and operates in the opposite dimension to the ConvNeXt block within each corresponding sub-operation of the query-key-value framework. Building upon the above two modules, we present a family of Fast-Forward Networks (FFNet). Our FFNet achieves remarkable performance improvements over previous state-of-the-art methods across a wide range of tasks. The strong and general performance of our proposed method validates our hypothesis and leads us to introduce “MetaMixer”, a general mixer architecture that does not specify sub-operations within the query-key-value framework. We show that using only simple operations like convolution and GELU in the MetaMixer can achieve superior performance. We hope that this intuition will catalyze a paradigm shift in the battle of network structures, sparking a wave of new research.
Figure 1: Overview of MetaMixer. (a) MetaMixer is derived by not specifying sub-operations within the query-key-value framework. We assert that the competence of Transformers primarily originates from MetaMixer, which we deem as the true backbone of Transformer. (b) To demonstrate this and propose a FFN-like efficient token mixer, we replace the inefficient sub-operations of self-attention with those from FFN while retaining MetaMixer structure. (c) Our MetaMixer-based pure ConvNets outperform domain-specialized competitors in various tasks, confirming the superiority of the MetaMixer framework.
@@ 1 Introduction
Transformer [94] models have emerged as the dominant backbone in Natural Language Processing (NLP) and have successfully extended their influence to the vision domains. The pioneering work of Vision Transformer (ViT) [27] and its follow-ups [7, 12, 59] have showcased promising performances in various vision tasks, posing a formidable challenge to the established Convolutional Neural Networks (CNNs) [34, 44]. Main component of the Transformer encoder, global self-attention, is specialized in modeling long-range dependencies but incurs a quadratic computational complexity with respect to the sequence length (resolutions). Therefore, deploying this module on resource-constrained mobile devices for real-time applications proves to be much more challenging than using lightweight CNNs. To address these issues, a lot of works have focused on modifying the self-attention module to be more vision-friendly [59, 65, 78, 32] or selectively applying it to low-resolution inputs [93, 111, 48, 107]. As such, effectively utilizing self-attention under strict latency constraints requires considerable efforts.
@@
On the other hand, most models effortlessly leverage Feed-Forward Networks (FFNs, also known as channel MLP), which aggregate channel information with moderate computational overhead. Nevertheless, in vision domains, FFN is usually excluded from analysis, although there are several motivations to consider FFNs. For example, regardless of modality, whether generative or perception tasks, FFNs serve as a fundamental building block in transformer-based architectures, such as GPT series, ViT and its diffusion-based variants [2, 69]; this ubiquity suggests that FFNs play an indispensable role in achieving competitive performance. Furthermore, Geva et al. [30] suggest that FFNs can be viewed as emulated neural key-value memories [82, 83], where the rows in the first projection weights act as keys correlating with input textual patterns, and the rows in the second projection weights serve as values that induce a distribution shift over the residual path. Based on this intuition, if a depthwise convolution is used as query projection, ConvNeXt block [60] can also be interpreted as operating via a query-key-value mechanism, similar to the self-attention (see Fig. 1). We thus hypothesize that the query-key-value framework, fundamental to the most successful mixers like self-attention and ConvNeXt block, is essential as the starting point for mixer design.
@@
We abstract the query-key-value mechanism into a generalized framework, MetaMixer, where sub-operations are not predefined, as illustrated in Fig. 1 (a). We further propose to replace the inefficient sub-operations of self-attention with those of FFN within the MetaMixer framework, aiming to bridge the efficiency gap between self-attention and FFN. We term it as FFNification. It consists of three aspects. (1) Key & Value Generation: Self-attention generates keys and values through linear projections (dynamic computation), whereas FFN employs static weights as keys and values. We thus adopt a static weight scheme for simple and efficient implementation. (2) Activation Function (Coefficient Generation): Attention softmax is costly because it involves calculating an exponent and summing across the token length, which poses challenges for parallelization [20]. Consequently, we replace the softmax with an element-wise activation function GELU [36]. (3) Query-Key-Value Interactions: The query-key and coefficient-value interactions in self-attention result in quadratic complexity $O(n^{2})$ with the token length $n$, as they compute relations between a single token and all other tokens. Instead, we explore whether these expensive calculations can be substituted by simple local aggregation like convolution. Inspired by large kernel CNNs [60, 26], we primarily use kernel sizes of 7 or even larger. The resulting FFNified attention and ConvNeXt block are symmetrical in terms of mixing dimension, and are thus adopted as the token mixer and channel mixer, respectively.
@@
In this paper, we validate the MetaMixer framework by conducting extensive experiments across various tasks leveraging only convolutions. Astonishingly, as shown in Fig. 1 (c), our MetaMixer-based methods outperform domain-specialized competitors, requiring only minor task-specific modifications. Based on these results, we argue that MetaMixer stands as the pivotal mixer architecture that ensures competitive performance. We make the following contributions:
@@
  • We provide an in-depth analysis of the query-key-value mechanism of FFN in vision models (Sec. 2 and Appendix A) and further abstract this mechanism into a general mixer architecture MetaMixer that encompasses the main modules of both Transformer and ConvNeXt models (Sec. 3.1).
@@
  • Although self-attention and FFN-based convolutional mixers can both be integrated within the MetaMixer, a significant efficiency gap remains between them. To address this, we propose a mixer design that adapts self-attention by employing FFN’s sub-operations and mimicking its functionality through the use of large kernel convolutions (Sec. 3.2). The resulting token mixer, termed FFNified attention, not only surpasses the performance of attention-based mixers but also achieves remarkable inference speedups.
  • Our methods achieve comprehensive state-of-the-art across a variety of modalities (image, point cloud, and time-series) and tasks (Sec. 4). We validate the efficacy of our approach on numerous tasks and comprehensively analyze the mixer design space through the MetaMixer framework (Appendix B, C).
@@
Figure 2: Key-Value Mechanism of FFNs in Visual Context. (a) Coefficient Sparsity: Astonishingly, the final stage shows significantly higher sparsity. To categorize numerous classes, a small subset (< 10%) of neurons are activated by the input. (b) Coefficient Map Corresponding to the Most Activated Key in the Last Layer: Keys specialized for each class selectively correlate at the target positions, suggesting their potential role in capturing distinctive visual features. The numbers in parentheses indicate the average values of the coefficients.
2 Feed-Forward Networks Are Key-Value Memories
@@
$Q=XW_{Q},K=XW_{K},V=XW_{V}$ (2.1)
$\mathrm{Attention}(X)=\mathrm{Softmax}(QK^{\top})V$ (2.2)
$\hskip 27.03003pt\mathrm{FFN}(X)=\mathrm{GELU}(XW_{1}^{\top})W_{2}$ (2.3)
@@
Recent NLP literatures [30, 66, 18, 98] also delve into the knowledge (or skills) encoded within FFNs through viewing them as key-value memories [83]. For instance, Geva et al. [30] find that each key correlates with a range of human-interpretable patterns, from shallow to more semantic ones; Dai et al. [18] and Meng et al. [66] update specific factual associations by identifying and editing coefficients that express certain factual knowledge. Also, from a different perspective, recent studies find that coefficient sparsity in FFNs (with ReLU [29] inserted) is a prevalent phenomenon [49, 117], and utilizing sparsity after ReLU significantly reduces the computational requirements of LLMs [67].
While analyses of FFNs have reached maturity in NLP, the computer vision community lacks similar investigations. we thus analyze the FFNs within PoolFormer [110] from a key-value memory perspective. The simplicity of pooling as the token mixer makes FFN the dominant computational component in terms of parameter count and capacity, highlighting its crucial role in the model’s overall performance. As shown in Fig. 2 (a), similar to the observations in [67, 49], layers in the final stage have < 10% activated coefficients. Considering that each value can be interpreted as a distribution over the output prediction for classification [30], the result is intuitive. Also, the coefficient map examples in Fig. 2 (b) demonstrate that class-specific keys consistently correlate with target objects. Further examples and a more detailed discussion are provided in Appendix A. These results demonstrate that FFNs are indeed point-wise key-value memories, detecting and aggregating input patterns through interpretable inner workings. Therefore, it is noteworthy that the two main modules of Transformer operate within the query-key-value framework.
3 Method
3.1 MetaMixer
@@
$c_{i}=\mathrm{Activation}(f(q,k_{i}))$ (3.1)
where $f(\cdot)$ is a compatibility function that computes the similarity between the query and the corresponding key (referred to as query-key interaction); $\mathrm{Activation}(\cdot)$ generates the coefficient $c$ on the values $v$. The output is then computed by an aggregation function $g(c,v)$ that sums the values, each weighted by the coefficient from Eq. 3.1. We refer to the process of aggregating values with their associated coefficients as coefficient-value interaction.
MetaMixer, as a general mixer architecture, can be instantiated as various well-established mixers by specifying sub-operations (query projection, key-value generation, query-key / coefficient-value interactions, and activation function). This nature demonstrates the capabilities of MetaMixer’s modular design. In Appendix B, we provide a detailed illustration of MetaMixer’s generality.
Figure 3: Overview of FFNification and Fast-Forward Network block. (a-b) Comparison between self-attention and FFNified attention; (c) Our mixer design easily adapts by selecting the convolution type based on the modality dimension.
@@ 3.2 FFNification
While in the Sec. 2, we have shown that both self-attention and FFN can be interpreted within the query-key-value mechanism, we note that self-attention (or local window-attention [59]) requires labor-intensive tuning and is less efficient in vision domains compared to FFN-based convolutional mixers [60, 85]. Hence, to incorporate the computational benefits of FFNs in diverse scenarios and further validate the significance of the MetaMixer beyond self-attention, we perform architectural surgeries that replace computationally intensive operations in self-attention with efficient alternatives. We refer to this process as FFNification. Specifically, we show that replacing the expensive pairwise dot-product in query-key/coefficient-value interactions with depthwise convolution is feasible, and that softmax is no longer an essential component.
@@
The most critical aspect of the FFNification process is leveraging depthwise convolution for query-key and coefficient-value interactions. In self-attention, as inter-token module, each query (token) interacts with all instance-specific key-value pairs through dot-product operations. In contrast, with depthwise convolution, a query (channel) interacts with static key-value (kernel) in a sliding window manner. FFNified attention, therefore, utilizes key-value pairs to compute local similarities and aggregate information (see Fig. 3 (a-b)). Leveraging convolutional interaction is not only efficient but also benefits from inductive biases such as locality and weight sharing across positions, which reduces model complexity and aids in learning translationally equivalent representations. To enlarge the receptive field, we mainly utilize a 7$\times$7 kernel size for image recognition tasks, while for other tasks, the kernel size is adjusted based on the characteristics of the modality or task. Moreover, to enhance performance without increasing inference cost, we leverage a structural re-parameterization [25], training an additional small kernel branch (e.g., 3$\times$3) alongside a large kernel, which are subsequently merged into the large kernel during inference (see Appendix F for a detailed description).
@@
While softmax self-attention has long been the de facto choice, its suitability in sparse computation scenarios like depthwise convolution warrants reconsideration. The exponential normalization in softmax tends to be highly centralized in a small number of values [61, 79, 5], which means convolution with a local receptive field may fail to achieve sufficient context aggregation. In contrast, GELU [36] do not suffer from this issue. Therefore, we adopt GELU in our settings.
@@
The derived mixer, FFNified attention, along with the ConvNeXt block [60], consists solely of hardware-friendly operations such as convolution and GELU. Furthermore, within the MetaMixer framework, the FFNified attention and ConvNeXt block are symmetrically positioned to operate in opposite dimensions across each sub-operation (see Fig. 1 (b)). Consequently, we employ FFNified attention and the ConvNeXt block (7$\times$7 depthwise convolution + FFN) as token mixer and channel mixer, respectively. This arrangement facilitates efficient and compact feature extraction.
3.3 Fast-Forward Network as a General Backbone

### v2

# FFNet: MetaMixer-based Efficient Convolutional Mixer Design

@@ Abstract
11footnotetext: Corresponding author.
Transformer, composed of self-attention and Feed-Forward Network (FFN), has revolutionized the landscape of network design across various vision tasks. While self-attention is extensively explored as a key factor in performance, FFN has received little attention. FFN is a versatile operator seamlessly integrated into nearly all AI models to effectively harness rich representations. Recent works also show that FFN functions like key-value memories. Thus, akin to the query-key-value mechanism within self-attention, FFN can be viewed as a memory network, where the input serves as query and the two projection weights operate as keys and values, respectively. Based on these observations, we hypothesize that the importance lies in query-key-value framework itself for competitive performance. To verify this, we propose converting self-attention into a more FFN-like efficient token mixer with only convolutions while retaining query-key-value framework, namely FFNification. Specifically, FFNification replaces query-key-value interactions with large kernel convolutions and adopts GELU activation function instead of softmax. The derived token mixer, FFNified attention, serves as key-value memories for detecting locally distributed spatial patterns, and operates in the opposite dimension to the ConvNeXt block within each corresponding sub-operation of the query-key-value framework. Building upon the above two modules, we present a family of Fast-Forward Networks (FFNet). Despite being composed of only simple operators, FFNet outperforms sophisticated and highly specialized methods in each domain, with notable efficiency gains. These results validate our hypothesis, leading us to propose “MetaMixer”, a general mixer architecture that does not specify sub-operations within the query-key-value framework. Building on the MetaMixer framework, we also introduce a hybrid strategy that harmoniously integrates attention and FFNified attention, offering a comprehensive view of mixer design.
@@ 1 Introduction
Transformer [95] models have emerged as the dominant backbone in Natural Language Processing (NLP) and have successfully extended their influence to the vision domains. The pioneering work of Vision Transformer (ViT) [27] and its follow-ups [7, 13, 58] have showcased promising performances in various vision tasks, posing a formidable challenge to the established Convolutional Neural Networks (CNNs) [37, 45]. Main component of the Transformer, global self-attention, is specialized in modeling long-range dependencies but incurs a quadratic computational complexity with respect to the sequence length. Therefore, deploying this module on resource-constrained mobile devices for real-time applications proves to be much more challenging than using lightweight CNNs. To address these issues, a lot of works have focused on modifying the self-attention module to be more vision-friendly [58, 64, 78, 33] or selectively applying it to low-resolution inputs [94, 110, 49, 106]. As such, effectively leveraging self-attention under strict latency constraints requires laborious adjustments.
@@
On the other hand, most models effortlessly leverage Feed-Forward Networks (FFNs, also known as channel MLP), which aggregate channel information with moderate computational overhead. Nevertheless, in vision domains, FFN is usually excluded from analysis, although there are several motivations to consider FFNs. For example, regardless of modality, whether generative or perception tasks, FFNs serve as a fundamental building block in transformer-based architectures, such as GPT series, ViT and its diffusion-based variants [2, 69]; this ubiquity suggests that FFNs play an indispensable role in achieving competitive performance. Furthermore, Geva et al. [30] suggest that FFNs can be viewed as emulated neural key-value memories [83, 82], where the rows in the first projection weights act as keys correlating with input patterns, and the rows in the second projection weights serve as values that induce a distribution shift over the residual path. Based on this intuition, if a depthwise convolution is used as query projection, ConvNeXt block [59] can also be interpreted as operating via a query-key-value mechanism, similar to the self-attention (see Fig. 1). We thus hypothesize that the query-key-value framework, fundamental to the most successful mixers like self-attention and ConvNeXt block, is essential as the starting point for mixer design.
Figure 1: Overview of MetaMixer. (a) MetaMixer is derived by not specifying sub-operations within the query-key-value framework. We assert that the competence of Transformers primarily originates from MetaMixer, which we deem as the true backbone of Transformer. (b) To demonstrate this and propose a FFN-like efficient token mixer, we replace the inefficient sub-operations of self-attention with those from FFN within MetaMixer structure.
@@
We thus abstract the query-key-value mechanism into a generalized framework, MetaMixer, where sub-operations are not predefined, as illustrated in Fig. 1 (a). We further propose to replace the inefficient sub-operations of self-attention with those of FFN within the MetaMixer framework, aiming to bridge the efficiency gap between self-attention and FFN. We term it as FFNification. It consists of three aspects. (1) Key-Value Generation: Self-attention generates keys and values through linear projections (dynamic computation), whereas FFN employs static weights as keys and values. We thus adopt a static weight scheme for simple and efficient implementation. (2) Activation Function (Coefficient Generation): Attention softmax is costly because it involves calculating an exponent and summing across the token length, which poses challenges for parallelization [20]. Consequently, we replace the softmax with an element-wise activation function GELU [38]. (3) Query-Key-Value Interactions: The query-key and coefficient-value interactions in self-attention result in quadratic complexity $O(n^{2})$ with the token length $n$, as they compute relations between a single token and all other tokens. Instead, we explore whether these expensive calculations can be efficiently substituted by simple local aggregation like convolution. Inspired by large kernel CNNs [59, 26], we primarily use kernel sizes of 7 or even larger. The resulting FFNified attention and ConvNeXt block are symmetrical in terms of mixing dimension (see Fig. 1 (b)), and are thus adopted as the token mixer and channel mixer, respectively.
@@
In this paper, we validate the MetaMixer framework by conducting extensive experiments across various tasks leveraging only convolutions. Our MetaMixer-based convolutional networks establish a strong and generally efficient baseline across a wide range of tasks and devices. Importantly, our convolutional mixer design, i.e., FFNified attention, surpasses highly specialized task-specific mixers with minimal modifications. Based on these results, we argue that MetaMixer stands as the pivotal mixer architecture that ensures competitive performance. We further discuss how to integrate existing mixer and FFNified attention within the MetaMixer framework, taking into account the characteristics of the task and the model’s data processing mechanism. We make the following contributions:
@@
  • We provide an in-depth analysis of the query-key-value mechanism of FFN in vision models and further abstract this mechanism into a general mixer architecture MetaMixer that encompasses the main modules of both Transformer and ConvNeXt models.
@@
  • Although self-attention and FFN-based convolutional mixers can both be integrated within the MetaMixer, a significant efficiency gap remains between them. To address this, we propose a mixer design that adapts self-attention by employing FFN’s sub-operations and mimicking its functionality through the use of large kernel convolutions. The resulting token mixer, termed FFNified attention, outperforms attention-based mixers in terms of the speed-accuracy trade-off, exhibiting robust efficiency across diverse devices.
  • Through a comprehensive comparison of our models with SOTA baselines across various modalities and tasks, we underscore the critical importance of the query-key-value framework (i.e., MetaMixer) and offer a strong and efficient baseline adaptable to various settings.
@@
Figure 2: Key-Value Mechanism of FFNs. (a) Coefficient Sparsity: Astonishingly, the final stage shows significantly higher sparsity. To categorize numerous classes, a small subset ($<$ 10%) of neurons are activated. (b) Coefficient Map corresponding to the most activated key in the last layer: Keys specialized for each class selectively correlate at the target regions, suggesting their potential role in capturing distinctive visual features. The numbers in parentheses indicate the average values of the coefficients.
2 FFNs Are Key-Value Memories
@@
$Q=XW_{Q},K=XW_{K},V=XW_{V},$ (1)
$\mathrm{Attention}(X)=\mathrm{Softmax}(QK^{\top})V,$ (2)
$\hskip 27.03003pt\mathrm{FFN}(X)=\mathrm{GELU}(XW_{1}^{\top})W_{2},$ (3)
@@
Recent NLP literatures [30, 65, 19, 98] also delve into the knowledge (or skills) encoded within FFNs through viewing them as key-value memories [82]. For instance, Geva et al. [30] find that each key correlates with a range of human-interpretable patterns, from shallow to more semantic ones; Dai et al. [19] and Meng et al. [65] update specific factual associations by identifying and editing coefficients that express certain factual knowledge. Also, from a different perspective, recent studies find that coefficient sparsity in FFNs is a prevalent phenomenon [51, 115].
While analyses of FFNs have reached maturity in NLP, the computer vision community lacks similar investigations. We thus analyze the FFNs within PoolFormer [109] from a key-value memory perspective. The simplicity of pooling as the token mixer makes FFN the dominant computational component in terms of parameter count and capacity, highlighting its crucial role in the model’s overall performance. As shown in Fig. 2 (a), consistent with the observations in [66, 51], layers in the final stage have $<$ 10% activated coefficients. Considering that each value can be interpreted as a distribution over the output prediction for each class [30], the result is intuitive. Also, the coefficient map examples in Fig. 2 (b) demonstrate that class-specific keys consistently correlate with target objects. Further examples and more detailed discussions are provided in Appendix A. These results demonstrate that FFNs are indeed point-wise key-value memories, detecting and aggregating input patterns through interpretable inner workings. Therefore, it is noteworthy that the two main modules of Transformer operate within the query-key-value framework.
3 Proposed Method
3.1 MetaMixer Framework
@@
$c_{i}=\mathrm{Activation}(f(q,k_{i})),$ (4)
where $f(\cdot)$ is a compatibility function that computes the similarity between the query and the corresponding key (referred to as query-key interaction); $\mathrm{Activation}(\cdot)$ generates the coefficient $c$ on the values $v$. The output is then computed by an aggregation function $g(c,v)$ that sums the values, each weighted by the coefficient from Eq. 4. We refer to the process of aggregating values with their associated coefficients as coefficient-value interaction.
Figure 3: Overview of FFNification and Fast-Forward Network block. (a-b) Comparison between self-attention and FFNified attention; (c) Our mixer design easily adapts by selecting the convolution type and kernel size based on the modality.
MetaMixer, as a general mixer architecture, can be instantiated as various well-established mixers by specifying sub-operations (query projection, key-value generation, query-key/coefficient-value interactions, and activation function). This nature demonstrates the capabilities of MetaMixer’s modular design. We provide a detailed illustration of MetaMixer’s generality in Appendix B.
@@ 3.2 FFNification
While in the Sec. 2, we have shown that both self-attention and FFN can be interpreted within the query-key-value framework, we note that self-attention (or window-attention [58]) requires labor-intensive tuning and is less efficient in vision domains compared to FFN-based convolutional mixers [59, 85]. Hence, to incorporate the computational benefits of FFNs in diverse scenarios and further validate the significance of the MetaMixer beyond self-attention, we perform architectural surgeries that replace computationally intensive operations in self-attention with efficient alternatives. We refer to this process as FFNification. Importantly, we show that replacing the expensive pairwise dot-product in query-key/coefficient-value interactions with depthwise convolution is feasible, and that softmax is no longer an indispensable component.
The most critical aspect of the FFNification process is employing depthwise convolution for query-key and coefficient-value interactions. In self-attention, as inter-token module, each query (token) interacts with all instance-specific key-value pairs through dot-product operations. In contrast, with depthwise convolution, a query (channel) interacts with static key-value (kernel) in a sliding window manner. FFNified attention, therefore, utilizes key-value pairs to compute local similarities and aggregate information (see Fig. 3 (a-b)).
@@
Leveraging convolutional interaction is not only efficient but also benefits from inductive biases such as locality and weight sharing across positions, which reduces model complexity and aids in learning translationally equivalent representations. To enlarge the receptive field, we mainly utilize a 7$\times$7 kernel size for image recognition tasks, while for other tasks, the kernel size is adjusted based on the characteristics of the modality and task. Moreover, to enhance performance without increasing inference cost, we leverage a structural re-parameterization [24], training an additional small kernel branch (e.g., 3$\times$3) alongside a large kernel, which are subsequently merged into the large kernel during inference (see Appendix E for a detailed description).
@@
While softmax self-attention has long been the de facto choice, its suitability in sparse computation like depthwise convolution warrants reconsideration. The exponential normalization in softmax tends to be highly centralized in a small number of values [61, 79, 5], which means convolution with a local receptive field may fail to achieve sufficient context propagation. In contrast, GELU do not suffer from this issue. Therefore, we adopt GELU in our settings.
@@
The derived mixer, FFNified attention, along with the ConvNeXt block [59], consists solely of hardware-friendly operations such as convolution and GELU. Furthermore, within the MetaMixer, the FFNified attention and ConvNeXt block are symmetrically positioned to operate in opposite dimensions across each sub-operation (see Fig. 1 (b)). Consequently, we employ FFNified attention and the ConvNeXt block (7$\times$7 depthwise convolution + FFN) as token mixer and channel mixer, respectively. This arrangement facilitates efficient and compact feature extraction.
3.3 FFNet as a General Backbone